1/1

James Yorkston,

Molly Malone

In Dublin’s fair city, Where girls are so pretty, I first set my eyes on sweet Molly Malone, As she pushed her wheelbarrow Through streets broad and narrow, Crying, Cockles and mussels, alive, alive oh! Alive, alive oh! alive, alive oh! Crying, Cockles and mussels, alive, alive oh!

She was a fishmonger, And sure twas no wonder, For so were her mother and father before, And they each wheeled their wheelbarrow, Through streets broad and narrow, Crying, Cockles and mussels, alive, alive oh! Alive, alive oh! alive, alive oh! Crying, Cockles and mussels, alive, alive oh!

She died of a fever, And no one could save her, And that was the end of sweet Molly Malone. Now her ghost wheels her barrow, Through streets broad and narrow, Crying, Cockles and mussels, alive, alive oh! Alive, alive oh! alive, alive oh! Crying, Cockles and mussels, alive, alive oh!

Verkinto de tiu ĉi Angla poemo estas James Yorkston.

Arg-374-733 (2013-11-25 12:31:14)

Pri la kanto vidu la retejojn http: // homepage. eircom. net/ ~seanjmurphy/ irhismys/ molly. htm kaj http: // homepage. eircom. net/ ~seanjmurphy/ irhismys/ mollylyrics. htm . En la unua retejo estas indikitaj ankaŭ la muziknotoj. Vidu ankaŭ la germanan version de la vikipediejo pri Molly Malone: http: // de. wikipedia. org/ wiki/ Molly_Malone . Legu ankaŭ: http: // en. wikisource. org/ wiki/ Molly_Malone .

http://www.poezio.net/version?poem-id=374&version-id=733